Night report 2017-03-24

Night report 2017-03-24

Observers: Francesca Onori, Aleksandra Hamanowicz

TAT on duty: Joe Anderson, Morgan Fraser

Support team: Morgan Fraser

Daytime calibrations: PESSTO_arcs, PESSTO_bias

Evening calibrations: none; strong wind

Weather: strong wind, can not open the dome untill UT 00:00. Thin clouds

Seeing: varying around 1.5'' (at UT 00:36); seeing 1.1'' (at UT 01:08); seeing 1.34'' (at UT 04:38), seeing 2.3'' (at UT 07.07)

UT/Target/RA/DEC/time spent(min)/slit/grism or filter, description

00:11 DES16C3cji 03:28:35.29 -27:09:06.5 44m 1.0'' gr13+r+i aborted after i images (problem with OB)

00:42 DES16C3cji 03:28:35.29 -27:09:06.5 40m 1.5'' gr13 aborted after 35min because too low, follow-up

01:20 LTT3864 10:32:13.90 -35:37:42.4 11m 1.0" + 1.5" Gr11 + Gr13 + Gr16 standard aborted due to wrong pointing

01:44 LTT3864 10:32:13.90 -35:37:42.4 11m 1.0" + 1.5" Gr11 + Gr13 + Gr16 standard

02:11 AT2017cfq 08:03:55.21 +26:31:12.6 50m 1.0'' gr11+gr16 internal lamp flat taken, follow-up

03:19 SN2017cbv 14:32:34.42 -44:08:02.7 20s 1.0'' gr11+gr16 internal lamp flat taken, follow up

03:35 AT2017awz 11:07:35.48 +22:51:04.5 50m 1.0'' gr11+gr16 internal lamp flat taken, follow up

04:36 Starting SOFI observations

04:58 AT2017awz 11:07:35.48 +22:51:04.5 06m JHK follow-up

05:08 AT2016jbu 07:36:25.96 -69:32:55.2 15m JHK aborted low quality image

05:19 image analysis

05:29 AT2016jbu 07:36:25.96 -69:32:55.2 15m JHK earthquake follow-up

05:47 SN2017cbv 14:32:34.42 -44:08:02.7 24m 1.0" BG+RG follow up

06:41 Hip070976 14:31:01.54 -43:50:53.40 10m 1.0" BG+RG Telluric

06:51 VVV2013 17:07:18.91 -39:06:26.4 23m JHK follow up

07:25 Starting EFOSC

08:02 LSQ15adm 20:14:21.82 -58:08:20.2 10m V seeing 1.2'' , follow up

08:16 AT2017cin 16:14:00.65 +26:54:58.6 15m 1.0'' gr13 classification

08:23 AT2017ceq 10:57:39.76 -41:58:33.7 5m 1.0" gr13 classification

09:00 AT2017cju 15:38:34.06 +13:20:52.4 10m 1.0'' gr13 classification (already classified)

09:24 AT2016cvk 22:19:49.39 -40:40:03.2 10m gri follow up

09:37 ATLAS17byq 17:07:47.34 +09:49:20.8 5m 1.5" gr13 classification